# Automated Bug Generation in the era of Large Language Models

### v1

@@ 5.2. RQ1: Quality of the Generated Bugs
  • # Statements involved in bug generation: This metric indicates the number of statements added, removed, or modified to generate the bugs. On average, even with the threshold values of $k=10\%$, BugFarm changes more statements to generate bugs compared to LEAM and $\mu$BERT. Although #SI is higher for BugFarm bugs, given that these statements are among the least attended statements, we will later see that the models will have a harder time distinguishing them from the correct code (more details in section 5.3).
@@
  • # Lines deleted: Deleting an entire statement, while likely to create a bug, may not be the best bug-inducing transformation, as it will change attention significantly and result in a runtime exception rather than a semantic bug. So, we also wanted to compare the generated bugs from this aspect. Columns LD in Table 1 show the percentage of bugs created with only lines deleted (or commented) during bug generation. BugFarm does not delete (or comment) any statement through bug-inducing transformation ($0\%$ on average for all the projects), compared to LEAM ($13\%$) and $\mu$BERT ($0.02$%).
@@
  • Edit distance: We also wanted to compare the edit distance between the original and buggy versions generated by each technique. We used Levenshtein (Levenshtein et al., 1966) edit distance, which measures the minimum number of single-character edits—insertions, deletions, or substitutions—required to change one string into another. Each code is represented as a string of characters to compute the Levenshtein edit distance. The results for edit distance normalized by #CB are available under columns ED in Table 1. Compared to LEAM and $\mu$BERT, BugFarm’s bugs have higher ED values. This again indicates that BugFarm’s changes to the code are bigger yet less noticeable, as we will show in RQ2 and RQ3. Between $\mu$BERT and LEAM, the latter generates bugs different from the original code in more places.
@@
  • Uniqueness: Finally, we were interested to see how much our bugs overlap with those generated by $\mu$BERT and LEAM. To that end, we measured the Exact Match (EM) and CodeBLEU (Ren et al., 2020) values between each BugFarm bug with all the corresponding bugs generated by $\mu$BERT and LEAM. For example, for a given method A, if BugFarm generates three bugs ${b_{1},b_{2},b_{3}}$ and LEAM generates four ${l_{1},l_{2},l_{3},l_{4}}$, we construct $12$ pairs of $\langle b_{i},l_{j}\rangle$ to compute the EM and CodeBLEU. EM is a strict all-or-nothing metric; being off by a single character results in a score of 00. If the characters of $b_{i}$ exactly match the characters of $b_{j}$, EM = $1$ for the pair; otherwise, EM = 00. CodeBLEU is a metric to measure weighted n-gram match between the pairs by considering not just the code tokens but also code syntax via abstract syntax trees (AST) and code semantics via data flow.
@@
    The EM and CodeBLUE values are shown under columns OL (overlap with LEAM) and OM (overlap with $\mu$BERT) in Table 1. The first number indicates EM, and the second one is CodeBLEU. These numbers show that only $7\%$ and $3.49\%$ of the total bugs generated by BugFarm for all the projects overlap with LEAM and $\mu$BERT, respectively, confirming the uniqueness of BugFarm bugs. The average CodeBLEU values are also $0.67$ and $0.75$. The high number for CodeBLEU is not a threat here, considering bug transformations do not greatly change the code compared to the original one. In fact, it shows that although BugFarm bugs are better at challenging learning-based techniques, they are semantically similar to LEAM and $\mu$BERT bugs; potentially as effective as them in their evaluated tasks (Tian et al., 2022).
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Compared to other techniques, BugFarm’s bug-inducing transformations involve more statement modification and change of several code tokens. The overlap between BugFarm bugs and other approaches is low, demonstrating their uniqueness.
5.3. RQ2: Effectiveness in Generating Hard-to-Detect Bugs
In this research question, we investigate the effectiveness of bug prediction models on BugFarm bugs compared to alternative approaches. Given the unavailability of off-the-shelf models as discussed in section 5.1, we fine-tuned three pre-trained code-language models (CodeBERT, CodeT5, and NatGen) using real-world and synthetic bugs. We adapted best practices for fine-tuning transformer models and used the same class distribution in fine-tuning, validating, and testing bug prediction models. All models were fine-tuned for at most $10$ epochs with loss-based early-stopping criteria of two consecutive epochs 555this is the default setting for fine-tuning CodeT5. To ensure a fair comparison, we checked that all the models converged before $10$ epochs. and selected the model with the least validation error in our experiments. For more reliable results, we repeated the evaluation 10 times and reported the average values.
Generally, one should run BugFarm on each fine-tuned model to generate bugs accordingly. However, our investigation showed that fine-tuning does not greatly impact the set of LAS (Algorithm 1 in section 4.2). This is consistent with the findings of prior work that shows fine-tuning only changes the attention of the last few layers, not impacting the overall attention of the model (Shi et al., 2023). It also shows that many of the generated BugFarm bugs in this study are reusable by other researchers. Consequently, we only generated bugs for the methods per each model whose LAS set was changed compared to the baseline pre-trained model.
5.3.1. Fine-tuning on synthetic bugs.
We fine-tuned the three baseline pre-trained models on LEAM and $\mu$BERT bugs. This provided us with six fine-tuned bug prediction models, namely, $\mu$BERT-CodeBERT, $\mu$BERT-CodeT5, $\mu$BERT-NatGen, LEAM-CodeBERT, LEAM-CodeT5, and LEAM-NatGen. Our goal is not only to evaluate these models in predicting BugFarm bugs, but also to see the extent to which BugFarm bugs are harder to be detected than other techniques. To that end, we evaluated the fine-tuned bug prediction models on BugFarm and the other approach, whose bugs were not used for fine-tuning. For example, we evaluated $\mu$BERT-CodeBERT on bugs generated by BugFarm and LEAM. To ensure a fair comparison, we considered methods per each project that both comparing techniques had generated confirmed bugs for them. This will result in the same FPR values for both techniques, but allows us to focus on the effectiveness of bugs for evaluating the model better. This is important from the security perspective since missing a bug/vulnerability is more severe than marking a correct code as buggy (Ami et al., 2023).
The rows 1–12 in Table 2 show the result of this experiment. For Accuracy, Precision, Recall, and F1 score, the lower metric value indicates the approach’s superiority (bugs are harder to distinguish from correct code, hence being detected). For FNR, a higher metric value indicates the bugs are harder to detect. From these results, we can clearly see that BugFarm bugs always achieve higher values for FNR (average margin of 23% (min=2.76%,max=40.89%)) and lower values for the other metrics (average margin of Accuracy=5.97%, Precision=1.48%, Recall=15.5%, and F1-score=10.78%). Furthermore, by looking at the F1 score values, we observe that the models fine-tuned on LEAM have a harder time detecting BugFarm bugs than those fine-tuned on $\mu$BERT. We believe this is because $\mu$BERT bugs are more diverse due to changing many tokens of the code and combining them through beam search, compared to LEAM bugs that try to mimic the bugs in their dataset, scrapped from the GitHub issue trackers.
5.3.2. Fine-tuning on real-world bugs.
To avoid any bias in our conclusion based on synthetic bugs, we also fine-tuned baseline pre-trained models with real-world bugs from three datasets, namely Defects4J (Just et al., 2014a) [^6] We could use the bugs from Mockito and Closure projects, which were excluded from our subjects since BugFarm only supports Maven at this time, BugSwarm (Tomassi et al., 2019), and RegMiner (Song et al., 2022). The real-world evaluation dataset consists of $723$, $3285$, and $36412$ original and buggy methods from Defects4J, BugSwarm, and RegMiner, respectively. This resulted in three bug prediction models, i.e., REAL-CodeBERT, REAL-CodeT5, and REAL-NatGen. We evaluated each model on BugFarm, $\mu$BERT, and LEAM bugs. The rows 13–21 in Table 2 show the results of this experiment, with margins indicating the difference with respect to second-best synthetic bug dataset. These results show similar trends we observed with models fine-tuned on synthetic bugs, i.e., BugFarm bugs result in higher FNR (54.35% on average) and lower Accuracy (50.86% on average), Precision (51.52% on average), Recall (45.64% on average), and F1 score (47.34% on average) values.
We can also see that models fine-tuned on real-world bugs underperform those fine-tuned on synthetic bugs across all metrics. A possible justification for this observation is the distribution shift between real-world bugs (used for fine-tuning) and synthetic bugs (used for testing). The two root causes for these distribution shifts are (1) the real-world bugs belonging to projects different than those from which we generated the synthetic bugs; and (2) the nature of real-world bugs is different from synthetic bugs. Especially for BugFarm and $\mu$BERT, the bug generation objective does not include any similarity to real-world bugs.
@@ 5.3. RQ2: Effectiveness in Generating Hard-to-Detect Bugs
Table 2. Effectiveness of studied fine-tuned models in predicting synthetic bugs. Each distinct set of rows show the same pre-trained model that is fine-tuned on the same dataset, but tested on different bug datasets.
Fine-tune-Model-Test Acc Prec Rec F1 FPR FNR
1 $\mu$BERT-CodeBERT-BugFarm 71.95 (4.40% normal-$\downarrow$) 90.04 (1.17% normal-$\downarrow$) 49.35 (11.83% normal-$\downarrow$) 63.76 (8.05% normal-$\downarrow$) 5.46 (0.00% --) 50.65 (15.04% normal-$\uparrow$)
2 $\mu$BERT-CodeBERT-LEAM 75.26 91.11 55.97 69.34 5.46 44.03
@@
4 LEAM-CodeBERT-$\mu$BERT 77.92 95.75 58.43 72.57 2.59 41.57
@@
6 $\mu$BERT-CodeT5-LEAM 76.08 87.9 60.48 71.66 8.33 39.52
@@
8 LEAM-CodeT5-$\mu$BERT 78.4 91.43 62.66 74.36 5.87 37.34
@@
10 $\mu$BERT-NatGen-LEAM 75.43 85.44 61.3 71.38 10.44 38.7
@@
12 LEAM-NatGen-$\mu$BERT 76.72 90.91 59.39 71.84 5.94 40.61
@@
14 REAL-CodeBERT-$\mu$BERT 55.22 58.4 36.31 44.78 25.87 63.69
15 REAL-CodeBERT-LEAM 57.68 61.44 41.23 49.35 25.87 58.77
@@
17 REAL-CodeT5-$\mu$BERT 51.5 51.57 49.42 50.47 46.42 50.58
18 REAL-CodeT5-LEAM 53.79 53.77 53.99 53.88 46.42 46.01
@@
20 REAL-NatGen-$\mu$BERT 50.85 50.71 61.16 55.45 59.45 38.84
21 REAL-NatGen-LEAM 53.82 53.02 67.1 59.23 59.45 32.9
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Bug prediction models, regardless of whether the fine-tuning dataset is from a synthetic or real-world bug dataset, have a harder time detecting BugFarm bugs than other synthetic bugs.
5.4. RQ3: Effectiveness in Generating Hard-to-Repair Bugs
Table 3. Effectiveness of FitRepair in repairing synthetic bugs. SI: Statements involved.
LEAM (Tian et al., 2022) $\mu$BERT (Khanfir et al., 2023) BugFarm-CodeBERT BugFarm-CodeT5 BugFarm-NatGen
Total Bugs (SI=1,SI=2,SI¿2) 100 (84,15,1) 100 (92,8,0) 100 (52,26,22) 100 (34,25,41) 100 (50,16,34)
@@ 5.4. RQ3: Effectiveness in Generating Hard-to-Repair Bugs
To further demonstrate the complexity of our bugs, we evaluate the ability of FitRepair in repairing BugFarm, $\mu$BERT, and LEAM bugs. The default configuration of FitRepair aims to generate $5000$ patches for a single bug and will terminate if this takes more than five hours. Given the size of our subject bugs and the non-negligible patch validation time, we reduced the maximum number of patches to 100100100100777This is the max allowance, and the tool often generates a lower number of patches (confirmed by the developers of FitRepair that this does not result in drastic performance degradation) and evaluated it on $100$ bugs sampled from each bug dataset. When sampling, we controlled for selecting bugs from the same methods for all approaches. To select the methods, we sorted them based on the descending order of method size—measured by the number of characters—and picked the top $100$. Our rationale is that the potential locations for bug injection increase when methods are longer. As a result, the likelihood of observing different bugs produced by each technique is higher (similar to the illustrative example in section 2). The collected bugs are from $11$ out of $15$ subject projects.
@@
The first row in Table 3 shows the distribution of generated bugs based on the number of statements involved in bug generation (#SI from RQ1). Most subject bugs from LEAM and $\mu$BERT differ in only one line with the original code, while BugFarm bugs are more diverse concerning this metric. Each bug dataset took FitRepair two to seven hours to generate $35$ patches, on average. We validated all the generated patches for each bug, and if one of the patches results in a green test suite, we ignore other patches and consider the bug repaired by FitRepair. The validation process of $500$ bugs took over $50$ hours. The second row of Table 3 shows the percentages of bugs that FitRepair successfully patched.
@@
FitRepair successfully repaired 34% and 49% of the LEAM and $\mu$BERT bugs, respectively. In contrast, it can only repair 27%, 22%, and 23% of BugFarm bugs generated for each pre-trained baseline. This is not surprising, as APR techniques are known to perform better in repairing bugs with SI=1. In fact, 88% of correct patches for $\mu$BERT and LEAM only differ with the corresponding bug in one line. This value is only 64% for BugFarm, which implies a higher complexity of BugFarm’s one-line bugs compared to alternative approaches. Our deeper investigation of the nature of bugs shows that LEAM and $\mu$BERT bugs mostly change the conditional/branch statements. Since repair templates of FitRepair are specifically designed to repair bugs in conditional/branch statements (logical expression in if statement), it can repair LEAM and $\mu$BERT bugs better. In contrast, BugFarm is more creative in introducing bug injection due to the power of LLMs in code synthesis, and it considers locations that the model attends less to for injecting the bugs. Consequently, it can challenge learning-based APR techniques crafted to repair known bug patterns. By looking at the other patches that fixed bugs with SI¿=2, we observed the same pattern, i.e., there was at least one statement changing the conditional statements.
{tcolorbox}

### v2

@@ V-B RQ1: Characteristics of the Generated Bugs
  • # Statements involved: This metric indicates the number of statements added, removed, or modified to generate the bugs. On average, even with the threshold values of $k=10\%$, BugFarm changes more statements to generate bugs compared to LEAM and $\mu$BERT. Although #SI is higher for BugFarm bugs, given that these statements are among the least attended statements, we will later see that the models will have a harder time distinguishing them from the correct code (more details in section V-C).
@@
  • # Lines deleted: Deleting an entire statement is very likely to change attention significantly and result in a runtime exception rather than a semantic bug. Columns LD show the percentage of bugs created by only deleting or commenting statements. BugFarm does not delete (or comment) any statement through bug-inducing transformation ($0\%$ on average for all the projects), compared to LEAM ($12.43\%$) and $\mu$BERT ($0.02$%).
@@
  • Edit distance: We also wanted to compare the edit distance between the original and generated bugs. We used Levenshtein [25] edit distance, which measures the minimum number of single-character edits—insertions, deletions, or substitutions—required to change one string into another. The results for edit distance normalized by #CB are available under columns ED. Compared to LEAM and $\mu$BERT, BugFarm’s bugs have higher ED values. This indicates that BugFarm’s changes to the code are bigger yet less noticeable, as we will show in RQ2 and RQ3.
@@
  • Uniqueness: We were interested to see how much our bugs overlap with those generated by $\mu$BERT and LEAM. To that end, we measured the Exact Match (EM) and CodeBLEU [33] values between each BugFarm bug with all the corresponding bugs generated by $\mu$BERT and LEAM. For a given method A, if BugFarm generates three bugs ${b_{1},b_{2},b_{3}}$ and LEAM generates four ${l_{1},l_{2},l_{3},l_{4}}$, we construct $12$ pairs of $\langle b_{i},l_{j}\rangle$ to compute the EM and CodeBLEU. EM is a strict all-or-nothing metric; being off by a single character results in a score of 00. If the characters of $b_{i}$ exactly match the characters of $b_{j}$, EM = $1$ for the pair; otherwise, EM = 00. CodeBLEU is a metric to measure weighted n-gram match between the pairs by considering not just the code tokens but also code syntax via abstract syntax trees (AST) and code semantics via data flow.
@@
    The $<$EM,CodeBLEU$>$ values are shown under columns OL (overlap with LEAM) and OM (overlap with $\mu$BERT). These numbers show that only $7.6\%$ and $4\%$ of the total bugs generated by BugFarm for all the projects overlap with LEAM and $\mu$BERT, respectively, confirming the uniqueness of BugFarm bugs. The high number for CodeBLEU ($0.7$ on average for both) shows that although BugFarm bugs are better at challenging learning-based techniques (section V-C-section V-D), they are semantically similar to LEAM and $\mu$BERT bugs; potentially as effective as them in their evaluated tasks [39].
@@
TABLE II: Effectiveness of studied fine-tuned models in predicting synthetic bugs. Each set of rows shows the same pre-trained model that is fine-tuned on the same dataset but tested on different bug datasets.
Fine-tune-Model-Test Acc Prec Rec F1 FNR
1 $\mu$BERT-CodeBERT-BugFarm 71.88 (4.49% $\downarrow$) 89.94 (1.11% $\downarrow$) 49.26 (12.18% $\downarrow$) 63.66 (8.26% $\downarrow$) 50.74 (15.55% $\uparrow$)
2 $\mu$BERT-CodeBERT-LEAM 75.26 90.95 56.09 69.39 43.91
@@
4 LEAM-CodeBERT-$\mu$BERT 77.79 95.46 58.35 72.43 41.65
@@
6 $\mu$BERT-CodeT5-LEAM 76.04 87.76 60.52 71.64 39.48
@@
8 LEAM-CodeT5-$\mu$BERT 78.23 90.97 62.68 74.22 37.32
@@
10 $\mu$BERT-NatGen-LEAM 75.39 85.32 61.33 71.36 38.67
@@
12 LEAM-NatGen-$\mu$BERT 76.61 90.28 59.63 71.82 40.37
@@
14 REAL-CodeBERT-$\mu$BERT 55.14 58.26 36.24 44.68 63.76
15 REAL-CodeBERT-LEAM 57.62 61.38 41.12 49.25 58.88
@@
17 REAL-CodeT5-$\mu$BERT 51.45 51.52 49.15 50.31 50.85
18 REAL-CodeT5-LEAM 53.81 53.79 54.05 53.92 45.95
@@
20 REAL-NatGen-$\mu$BERT 50.81 50.67 61.33 55.49 38.67
21 REAL-NatGen-LEAM 53.81 53.01 67.05 59.21 32.9
Summary. Compared to other techniques, BugFarm’s bug-inducing transformations involve more statement modification and change of several code tokens. The overlap between BugFarm bugs and other approaches is low, demonstrating their uniqueness.
V-C RQ2: Effectiveness in Generating Hard-to-Detect Bugs
In this research question, we investigate the effectiveness of bug prediction models on BugFarm bugs compared to alternative approaches. Given the unavailability of off-the-shelf models as discussed in section V-A, we fine-tuned three pre-trained code-language models (CodeBERT, CodeT5, and NatGen) using real-world and synthetic bugs. We adapted best practices for fine-tuning transformer models and used the same class distribution in fine-tuning, validating, and testing bug prediction models. All models were fine-tuned for at most $10$ epochs with loss-based early-stopping criteria of two consecutive epochs444this is the default setting for fine-tuning CodeT5. To ensure a fair comparison, we checked that all the models converged before $10$ epochs.. We selected the model with the least validation error, repeated the evaluation $10$ times, and reported the average values.
Generally, one should run BugFarm on each model for bug generation. However, our investigation showed that fine-tuning does not greatly impact the set of LAS (Algorithm 1 in section IV-B). This is consistent with the findings of prior work that shows fine-tuning only changes the attention of the last few layers, not impacting the overall attention of the model [36]. It also shows that many of the generated BugFarm bugs in this study are reusable by other researchers. Consequently, we only generated bugs for the methods per each model whose LAS set was changed compared to the baseline pre-trained model.
V-C1 Fine-tuning on synthetic bugs
It is possible that a learning-based bug prediction technique uses synthetic bugs for training or fine-tuning. So, we first investigate how challenging BugFarm’s bugs are, compared to other synthetic bugs, to challenge such models. To that end, we fine-tuned the three pre-trained models on LEAM and $\mu$BERT bugs. This provided us with six fine-tuned bug prediction models, namely, $\mu$BERT-CodeBERT, $\mu$BERT-CodeT5, $\mu$BERT-NatGen, LEAM-CodeBERT, LEAM-CodeT5, and LEAM-NatGen. We then evaluated each on BugFarm and the other approach, whose bugs were not used for fine-tuning. For example, we evaluated $\mu$BERT-CodeBERT on bugs generated by BugFarm and LEAM. This pairwise comparison will show us which techniques generate bugs that are harder to detect.
The rows 1–12 in Table II show the result of this experiment. For Accuracy, Precision, Recall, and F1 score, the lower metric value indicates the approach’s superiority (bugs are harder to distinguish from correct code, hence being detected). For FNR, a higher metric value indicates higher FN, showing the model struggles more to detect them. From these results, we can clearly see that BugFarm bugs always achieve higher values for FNR (average margin of 23.31% (min=3.34%, max=40.53%)) and lower values for the other metrics (average margin of Accuracy=5.95%, Precision=1.2%, Recall=15.73%, and F1-score=10.84%).
Furthermore, F1 score values suggest the models fine-tuned on LEAM have a harder time detecting BugFarm bugs than those fine-tuned on $\mu$BERT. We believe this is because $\mu$BERT bugs are more diverse due to changing many tokens of the code and combining them through beam search, compared to LEAM bugs that try to mimic the bugs scrapped from the GitHub issue trackers.
TABLE III: Effectiveness of FitRepair in repairing synthetic bugs. SI: Statements involved.
LEAM [39] $\mu$BERT [24] BugFarm-CodeBERT BugFarm-CodeT5 BugFarm-NatGen
Total Bugs (SI=1,SI=2,SI$>$2) 200 (174,25,1) 200 (187,13,0) 200 (125,40,35) 200 (92,43,65) 200 (104,30,66)
@@ V-C1 Fine-tuning on synthetic bugs
Note that we have not used BugFarm bugs for fine-tuning models. The reason is that BugFarm bugs are in-distribution samples, and alterations are within the decision boundaries for models: we only accept mutants that do not change the model’s attention so that the code representation of buggy and correct code is similar (Algorithm 2 Lines 11–14). Fine-tuning on such examples theoretically cannot improve in-distribution performance and may even worsen out-of-distribution performance . As we claimed before, BugFarm bugs are for evaluating and challenging learning-based bug-related tasks and should be used for that purpose only.
V-C2 Fine-tuning on real-world bugs
To avoid any bias in our conclusion based on synthetic bugs, we also fine-tuned baseline models with real-world bugs from three datasets, namely Defects4J [23] [^5] We could use the bugs from Mockito and Closure projects, which were excluded from our subjects since BugFarm only supports Maven at this time, BugSwarm [40], and RegMiner [37]. The real-world evaluation dataset consists of $723$, $3285$, and $36412$ original and buggy methods from Defects4J, BugSwarm, and RegMiner, respectively. This results in three bug prediction models, i.e., REAL-CodeBERT, REAL-CodeT5, and REAL-NatGen. We evaluated each model on BugFarm, $\mu$BERT, and LEAM bugs. The rows 13–21 in Table II show the results of this experiment, with margins indicating the difference with respect to second-best synthetic bug dataset. These results show similar trends we observed with models fine-tuned on synthetic bugs, i.e., BugFarm bugs result in higher FNR (54.12% on average) and lower Accuracy (50.93% on average), Precision (51.59% on average), Recall (45.88% on average), and F1 score (47.5% on average) values.
That models fine-tuned on real-world bugs underperform those fine-tuned on synthetic bugs across all metrics. A possible justification for this observation is the distribution shift between real-world bugs (used for fine-tuning) and synthetic bugs (used for testing). The two root causes for these distribution shifts are (1) the real-world bugs belonging to projects different than those from which we generated the synthetic bugs; and (2) the nature of real-world bugs is different from synthetic bugs. Especially for BugFarm and $\mu$BERT, the bug generation objective does not include any similarity to real-world bugs.
Summary. Bug prediction models, regardless of whether the fine-tuning dataset is from a synthetic or real-world bug dataset, have a harder time detecting BugFarm bugs than other synthetic bugs.
V-D RQ3: Effectiveness in Generating Hard-to-Repair Bugs
@@
To further demonstrate the complexity of our bugs, we evaluate the ability of FitRepair in repairing BugFarm, $\mu$BERT, and LEAM bugs. We configured FitRepair to generate $100$ patches  [^6] The tool often generates fewer patches than this max allowance value. per bug and terminate if this takes more than five hours  [^7] Confirmed by FitRepair authors to ensure proper performance.. Given the time-consuming nature of program repair, we evaluated FitRepair on $1000$ generated bugs  [^8] This number is 2.5 times than the bugs in FitRepair evaluation., constitutes from $200$ sampled confirmed bugs from each bug dataset (LEAM, $\mu$BERT, BugFarm-CodeBERT, BugFarm-CodeT5, and BugFarm-NatGen). When sampling, we controlled the selection of bugs from the same methods for all approaches: we sorted them based on the descending order of method size—measured by the number of characters—and picked the top $200$. Our rationale is that the potential locations for bug injection increase when methods are longer. As a result, the likelihood of observing different bugs produced by each technique is higher (similar to the illustrative example in section II). The collected bugs are from $14$ out of $15$ subject projects.
@@
The first row in Table III shows the distribution of generated bugs based on the number of statements involved in bug generation (#SI from RQ1). Most subject bugs from LEAM and $\mu$BERT differ in only one line with the original code, while BugFarm bugs are more diverse concerning this metric. Each bug dataset took FitRepair two to seven hours to generate $35$ patches, on average. We validated all the generated patches for each bug, and if one of the patches results in a green test suite, we ignore other patches and consider the bug repaired by FitRepair. The validation process of $1000$ bugs took over $100$ hours. The second row of Table III shows the percentages of bugs that FitRepair successfully patched.
@@
FitRepair successfully repaired 36% and 49% of the LEAM and $\mu$BERT bugs, respectively. In contrast, it can only repair 29%, 29%, and 28% of BugFarm bugs generated for each baseline model. This is not surprising, as APR techniques are known to perform better in repairing bugs with SI=1. In fact, 87% of correct patches for $\mu$BERT and LEAM only differ with the corresponding bug in one line. This value is only 62% for BugFarm, which implies a higher complexity of BugFarm’s one-line bugs compared to alternative approaches.
Our deeper investigation of the nature of bugs shows that LEAM and $\mu$BERT bugs mostly change the conditional/branch statements. Since repair templates of FitRepair are specifically designed to repair bugs in conditional/branch statements (logical expression in if statement), it can repair LEAM and $\mu$BERT bugs better. In contrast, BugFarm is more creative in introducing bug injection due to the power of LLMs in code synthesis, and it considers locations that the model attends less to for injecting the bugs. Consequently, it can challenge learning-based APR techniques crafted to repair known bug patterns. By looking at the other patches that fixed bugs with SI>>>=2, we observed the same pattern, i.e., there was at least one statement changing the conditional statements.